Two samples are available. one is 24% protein and the other is 12% protein. how much of each sample should be used to get 900 pounds that is 16%?
Answer by Boreal(15235)

One way is that 16% is one-third of the way from 12% to 24%, so use 2/3 12% (600 pounds) and 1/3 24% (300 pounds)
x=12%
900-x=24%
.12x+.24(900-x)=900*.16
.12x+216-.24x=144
-.12x=-72
x=600 pounds
900-x=300 pounds